OpenStudy (anonymous):

Find the best possible bounds fo rthe following equation: y=e^-(x)^2 for absolute x which is less than or equal to .3

OpenStudy (anonymous):

it is .3

OpenStudy (turingtest):

\[y=e^{-x^2} ,\left| x \right| \le0.3\]

OpenStudy (anonymous):

ya

OpenStudy (turingtest):

I'm not sure what you mean by best possible bounds

OpenStudy (anonymous):

do u know abt lower bounds and upper bounds?

OpenStudy (anonymous):

cus i don;t!!!!

OpenStudy (anonymous):

I think it is in other words the yvalue of the global min and max

OpenStudy (anonymous):

upperbound is global max and lower bound is global min

OpenStudy (turingtest):

Right that is pretty much my understanding.

OpenStudy (anonymous):

i guess you would find the critical points and the end points which is .3

OpenStudy (turingtest):

Right, so we differentiate the above to get\[{dy \over dx}=-2xe^{-x^2}=0\to x=\left\{ 0,\infty \right\}\]only three points to check then, our critical point and the endpoints: f(-0.3) f(0) and f(0.3) f(-0.3)=e^0.9=2.46 f(0)=1 f(0.3)=e^-0.9=0.41 so I guess the bounds are those numbers...

OpenStudy (turingtest):

2.46 and 0.41 I mean

OpenStudy (anonymous):

sorry i was gone i just came back

OpenStudy (turingtest):

I dunno if this is what you want, but it's all I can think to do. I still don't know what it means by "best bounds" though.

OpenStudy (anonymous):

i mean the global min and global max

OpenStudy (turingtest):

well these are the global max and min within those bounds, so there you have it

OpenStudy (anonymous):

there shld not be a -.3

OpenStudy (anonymous):

since it is absolute x

OpenStudy (anonymous):

I have a question